OpenStudy (anonymous):

Chelsea invested $15,000 at 4% interest compounded annually. How much interest will she earn in 4 years?

OpenStudy (wolf1728):

Total = Principal * (1+rate)^years Total = $15,000 * (1.04)^4 Total = $15,000 * 1.16985856 Total = 17,547.88 Interest = 17,547.88 -15,000.00 Interest = 2,547.88 Calculator to check the answer http://1728.org/compint.htm Formulas: http://1728.org/compint2.htm